# Multimodal Conditional 3D Face Geometry Generation

### v1

@@ Abstract.
We present a new method for multimodal conditional 3D face geometry generation that allows user-friendly control over the output identity and expression via a number of different conditioning signals. Within a single model, we demonstrate 3D faces generated from artistic sketches, 2D face landmarks, Canny edges, FLAME face model parameters, portrait photos, or text prompts. Our approach is based on a diffusion process that generates 3D geometry in a 2D parameterized UV domain. Geometry generation passes each conditioning signal through a set of cross-attention layers (IP-Adapter), one set for each user-defined conditioning signal. The result is an easy-to-use 3D face generation tool that produces high resolution geometry with fine-grain user control.
1. Introduction
@@
The creation of 3D facial geometry for digital human characters is a modeling task that usually requires tremendous artistic skill. Digital sculpting with 3D modeling tools is a time-consuming and demanding process, especially when the target is as recognizable as a human face. This complexity has prompted research into data-driven sculpting methods (Gruber et al., 2020) and other, more user-friendly, interactive interfaces (Kim et al., 2015).
@@
Several common morphable 3D face models (e.g. FLAME (Li et al., 2017)) simplify the facial modeling task by providing a shape subspace to operate in, as well as simple parameters to control the identity and expression geometry without the need for 3D modeling skills, but they are limited in expressiveness and offer only basic control knobs.
@@
Recent methods can create high quality 3D geometry and textures from text prompts (Zhang et al., 2023a; Wu et al., 2023) via optimization, leveraging large pre-trained text-to-image diffusion models (Rombach et al., 2022). These methods allow layman users to create 3D faces through natural text descriptions. While this is a powerful approach, it can still be difficult to achieve a particular output through text description (Koley et al., 2024). Some concepts like the specific curvature of a face or a unique facial expression are much easier to convey via sketches, edge contours or portrait photos than through text.
@@
In the image domain, approaches like ControlNet (Zhang et al., 2023b) or T2I-Adapter (Mou et al., 2023) have demonstrated controllable image generation beyond text using sketches, images, or edge maps as conditioning signals. These methods provide users with much more fine-grained control over the generation process than text-based methods alone. Ye et al. (2023) propose IP-Adapters to control Stable Diffusion (Rombach et al., 2022) with image prompts by learning new cross-attention layers. However, image-based methods are not easy to extend to 3D facial geometry generation.
@@
We present a flexible new method for 3D facial geometry generation that creates high quality faces from any one of various inputs, including sketches, 2D landmarks, Canny edges, portrait photos, FLAME face model parameters and text. Our approach is to train a conditional diffusion model on a high quality 3D facial dataset constructed from high resolution scans (Chandran et al., 2020) represented in the 2D UV domain. Our model is trained from scratch, without the need for a pre-trained foundation model like Stable Diffusion. To condition our model we train one set of cross-attention layers for each type of conditioning input, following IP-Adapter (2023). First, the diffusion model learns to inject FLAME parameters via the original UNet cross-attention layers. We then freeze the diffusion model while training additional sets of cross-attention layers (e.g. one for artist sketches, 2D landmarks, portrait photos, etc). An interesting side effect of the FLAME-conditioned model is that it allows us to re-interpret FLAME-parameterized faces in a generative sense - providing a space of high resolution stochastic variations on top of the traditional low resolution FLAME model.
@@
Our method allows for fast and user-friendly creation of 3D digital character faces, controlled by the input mode preferred by the user, all within a single model. We demonstrate several applications of our method including sketch-based 3D face modeling, geometry from 2D facial landmarks, Canny edges, or portrait photos, text-to-3D facial geometry, and finally, extending the FLAME model space by allowing stochastic diffusion sampling conditioned on the same semantic parameters.
2. Related Work
@@
2.1. 3D Face Geometry Generation
@@
Recent work uses diffusion models to control the generation of novel 3D face geometry. ShapeFusion (Potamias et al., 2024) generates face geometry by running the diffusion process directly on the mesh input vertices. It allows unconditional and conditional face geometry generation and supports various editing operations on a given mesh, based on selected vertices (anchor points). However, it does not support conditioning signals beyond vertices. 4DFM (Zou et al., 2024) trains an unconditional diffusion model on a set of sparse 3D landmarks for facial expression generation. It can generate dynamic facial expression sequences based on 3D landmarks by retargeting the landmarks to a mesh after the diffusion process. While they support conditioning with different signals such as expression labels and text, they achieve control via classifier-guidance (Dhariwal and Nichol, 2021) which requires training additional classifiers on noisy data.
@@
Other methods focus on 3D face or head avatars, which can generate 3D geometry and texture. Rodin (Wang et al., 2023) can generate triplane-based head geometry with text or image conditioning, but the resulting geometry is extracted with Marching Cubes (Lorensen and Cline, 1987) and thus not in the same topology across generations. DreamFace (Zhang et al., 2023a), Wu et al. (2023) and Bergman et al. (2023) propose pipelines that generate 3D face geometry and textures from text. The geometry is created by optimizing 3D Morphable Model (3DMM) (Blanz and Vetter, 1999) parameters using a Score Distillation Loss (SDS) (Poole et al., 2023). During the optimization, the SDS loss uses the feedback from a pre-trained text-to-image latent diffusion model to update the 3DMM parameters given a geometry render. This setup requires a differentiable renderer and often generates over-smoothed results with little details when compared to diffusion sampling (Zhou et al., 2024). Additionally, they require Stable Diffusion (SD) (Rombach et al., 2022) which was pre-trained on billions of images (Schuhmann et al., 2022) to guide their generations.
@@
In our work, we generate controllable 3D face geometry in a single common topology from several different conditioning modes without reliance on SDS optimization, classifier-guidance or a SD prior.
2.2. Multimodal Conditional Image Generation
@@
Image generation with diffusion models can be controlled with conditioning modes that are different from text such as sketches (Voynov et al., 2023), Canny edges (Canny, 1986), RGB images (Ye et al., 2023), expression parameters (Kirschstein et al., 2024) or face shape (Ding et al., 2023; Gu et al., 2024). To control pre-trained diffusion models with new modes, ControlNet (Zhang et al., 2023b) introduces a trainable copy of the diffusion model’s UNet encoding blocks, which take the new conditioning as input. The output of the copied model is added to the skip-connections of the frozen pre-trained diffusion model. A separate trainable copy of the UNet encoding blocks (361M parameters) is created per conditioning mode. T2I-Adapter (Mou et al., 2023) aligns the internal knowledge of a pre-trained diffusion model with new control modes by proposing a small adapter network that achieves control similar to ControlNet, while requiring less parameters (77M). IP-Adapter (Ye et al., 2023) injects each conditioning via separate cross-attention layers (Vaswani et al., 2017) while requiring even less parameters (22M). It introduces new cross-attention layers whose outputs are added to those of the original UNet. Ye et al. (2023) show that the diffusion model follows the added conditioning signal closely, when it comes through the newly trained cross-attention layers. In general, adapters can add control with new conditioning modes even long after the training of the underlying base diffusion model is concluded. They can also add new conditioning modes for which only limited paired training data is available, because the underlying diffusion model is frozen, avoiding issues such as catastrophic forgetting (Kirkpatrick et al., 2017; Zhang et al., 2023b). While pre-trained text-to-image diffusion models understand the RGB image domain and can generate images conditioned on a large variety of input modalities, they are not trivially extendable to generating 3D face geometry instead. We represent 3D face geometry in the 2D UV domain, which enables us to train a 2D diffusion model that can incorporate new conditioning modes using IP-Adapters. However, even in the 2D UV domain, 3D geometry is still far from what pre-trained text-to-image diffusion models have seen before. Thus, we train our diffusion model from scratch.
3. Multimodal 3D Face Geometry Generation
@@
3.1. Dataset and Geometry Representation
@@
To represent our face geometries, we generate a novel dataset based on the 3D scan data acquired by Chandran et al. (2020), where all faces are stabilized and in topological correspondence. In total, we use 323 identities in our training dataset, where each identity shows 24 different facial expressions (7752 examples). We subtract a template face shape from all faces in the dataset and thus represent each individual face as a delta from the template face. The computed delta representation reduces artifacts in the generated 3D face geometry, when compared to the full face representation (Fig. 3). We transform each delta face into a vertex delta position map in UV space, which is suitable for being processed by neural networks (Feng et al., 2018; Otto et al., 2022). This representation records the x, y, z coordinates of the face geometry within a 3-channel image, similar to traditional color texture maps, but instead of an RGB value at each pixel we store the x, y, z delta values. To improve generalization, we augment our existing geometry training data by synthetic identities which we generate via identity interpolation (50k examples) and by mixing face parts of different identities together (Funkhouser et al., 2004) (150k examples). Adding the augmentations during LDM training improves the generalization to novel identities when conditioning on the FLAME parameters (Table 1). We use the parameter space of a common 3D morphable face model (FLAME (Li et al., 2017)) as a base conditioning because we can fit FLAME to the scan data and to the augmented data and thus generate a large dataset of paired geometry-FLAME parameter data. Additionally, we create paired training data for several conditioning modes that only have limited paired data available. For example, portrait photos are only available for the scans from the dataset of Chandran et al. (2020), but not for the augmentation data. However, it is possible to inject new modalities with limited paired data by training new cross attention layers while keeping the LDM frozen as shown by Ye et al.(2023) (and described in Section 3.4).
3.2. Variational Autoencoder
@@
To reduce the computational requirements for the diffusion model, we downsample our UV position map data by a factor of four into a latent space using a variational autoencoder (VAE) (Kingma and Welling, 2014) consisting of an encoder and a decoder . We train the VAE from scratch following the autoencoder loss function and architecture as it is presented in related work (Rombach et al., 2022; Esser et al., 2021). Specifically, we use the VQ-GAN (Esser et al., 2021) autoencoder loss:
@@
consists of a pixel-wise L1 loss and a LPIPS (Zhang et al., 2018) perceptual loss. It compares the input UV position maps to the reconstructions through the VAE. evaluates inputs and reconstructions with a patch-based discriminator (Isola et al., 2017) and employs a codebook loss which serves as a latent space regularizer. Please refer to Esser et al. (2021) for more details.
3.3. Latent Diffusion Model
@@
Next, we train a latent diffusion model (LDM) (Rombach et al., 2022), that learns to generate latent UV position maps . To train the LDM, a forward diffusion process is defined as a Markov chain, which noises the latents following a fixed noise schedule of uniformly sampled timesteps. At the last time step , the distribution is Gaussian. We can directly sample at an arbitrary timestep by:
@@
We learn to predict the noise that was added to a noisy latent image following Ho et al. (2020):
where is the timestep, is a FLAME parameter conditioning and is the UNet (Ronneberger et al., 2015) neural network with parameters . The FLAME conditioning is obtained by fitting FLAME to the face geometry encoded in and mapping it through a MLP .
During inference (reverse diffusion process) we generate latent 2D UV position maps from the model distribution. We start from and iteratively compute less noisy latents until we reach a clean latent sample . Sampling following DDPM (Ho et al., 2020) or DDIM (Song et al., 2021) computes from based on the UNet output.
3.4. Multimodal Conditional Generation
To control the generations with additional conditioning modes (beyond the FLAME parameters ), we train different sets of cross-attention layers, following IP-Adapter (Ye et al., 2023). The LDM itself is kept frozen. In this way, we can integrate novel conditioning modes post-LDM training even with limited paired mode-geometry data. We train one set for each of the following conditioning modes: sketches , Canny edges , 2D landmarks , and portrait photos . The output of the new cross-attention layers is added to the outputs of the existing LDM cross-attention layers, thereby injecting the new conditioning signal into the generation process:
@@ 3.4. Multimodal Conditional Generation
Prior to passing each of the above-mentioned conditioning modes to the cross-attention layers, we pass each of them through CLIP (Radford et al., 2021) and extract a 768 dimensional global CLIP feature vector, which serves as our conditioning representation. Following IP-Adapter, we train a small projection network consisting of one linear layer and layer normalization, designed to project the CLIP feature vector into several extra context tokens before injecting it into the cross-attention layers. We use 16 tokens for each of our conditionings to allow for meaningful attention computation.
@@
At inference time, we can control the 3D face geometry generation with any of the modes using the respective set of cross-attention layers and classifier-free guidance (Ho and Salimans, 2021). The strength of the conditioning signal can be increased by increasing the hyperparameter :
@@
To condition only on a newly added mode or to generate geometry unconditionally, the FLAME conditioning is set to its null embedding. Additionally, for unconditional generation, the new cross-attention in Eq. 4, which feeds to the diffusion model, is simply not added to the original attention. Our full method pipeline is visualized in Fig. 2.
3.5. Implementation Details
For our dataset, we crop the full head face geometry to allocate more vertices to the face region, representing 50520 face vertices within each UV position map. At resolution we can represent reasonably high-resolution face geometry while being able to limit our VAE training time to 8 days using our training dataset of 7752 samples (1.4 seconds/iteration; batch size 8). We use a learning rate of 4.5e-6 and a codebook size of 8192. Note that for training our VAE, we did not use the data augmentations described in Section 3.1 as the autoencoder was already able to reconstruct test geometries with high accuracy when trained only on the studio dataset. Next, we train our LDM for 4 days (1.6 seconds/iteration; batch size 12) with a learning rate of 1e-4 and diffusion timesteps . We utilize geometry data augmentations with corresponding FLAME fits during training (+200k samples) to allow for better generalization across identities during generation. Afterwards, we train each set of cross-attention layers with a learning rate of 1e-4 for 6 days (3.3 seconds/iteration; batch size 24) while keeping the LDM frozen. With a probability of 0.05, we randomly set either or or both to their null embeddings during training. This step enables classifier-free guidance at inference. Note that we do not add augmented geometry data to train the new cross-attention layers because we do not have access to paired mode-geometry data for modes such as portrait photos. All experiments were run on a single RTX A6000 GPU. Also note that the VAE and the LDM have to be trained only once and novel conditioning modes can be added by training only the new cross-attention layers. In the next section, we show and evaluate our results either on a separate held-out validation set which contains 18 different identities displaying 24 expression each (432 samples) or on unconstrained in-the-wild test data. Unless mentioned otherwise, we generate every result by running DDPM sampling steps with conditioning strength . In total, generating a geometry sample takes 11 seconds.
4. Results

### v2

@@ Abstract
We present a new method for multimodal conditional 3D face geometry generation that allows user-friendly control over the output identity and expression via a number of different conditioning signals. Within a single model, we demonstrate 3D faces generated from artistic sketches, portrait photos, Canny edges, FLAME face model parameters, 2D face landmarks, or text prompts. Our approach is based on a diffusion process that generates 3D geometry in a 2D parameterized UV domain. Geometry generation passes each conditioning signal through a set of cross-attention layers (IP-Adapter), one set for each user-defined conditioning signal. The result is an easy-to-use 3D face generation tool that produces topology-consistent, high-quality geometry with fine-grain user control.
keywords:
[…]
@@
The creation of 3D facial geometry for digital human characters is a modeling task that usually requires tremendous artistic skill. Digital sculpting with 3D modeling tools is a time-consuming and demanding process, especially when the target is as recognizable as a human face. This complexity has prompted research into data-driven sculpting methods [1] and other, more user-friendly, interactive interfaces [2].
@@
Several common morphable 3D face models (e.g. FLAME [3]) simplify the facial modeling task by providing a shape subspace to operate in, as well as simple parameters to control the identity and expression geometry without the need for 3D modeling skills, but they are limited in expressiveness and offer only basic control knobs.
@@
Recent methods can create high quality 3D geometry and textures from text prompts [4, 5, 6, 7, 8] via optimization, leveraging large pre-trained text-to-image diffusion models [9]. These methods allow layman users to create 3D faces through natural text descriptions. While this is a powerful approach, it can still be difficult to achieve a particular output through text description [10]. Some concepts like the specific curvature of a face or a unique facial expression are much easier to convey via sketches, edge contours or portrait photos than through text.
@@
In the image domain, approaches like ControlNet [11] or T2I-Adapter [12] have demonstrated controllable image generation beyond text using sketches, images, or edge maps as conditioning signals. These methods provide users with much more fine-grained control over the generation process than text-based methods alone. Ye et al. [13] propose IP-Adapters to control Stable Diffusion [9] with image prompts by learning new cross-attention layers. However, image-based methods are not easy to extend to 3D facial geometry generation.
@@
We present a flexible new method for 3D facial geometry generation that creates high quality faces from any one of various inputs, including sketches, 2D landmarks, Canny edges, FLAME face model parameters, portrait photos and text. Our approach is to train a conditional diffusion model on a high quality 3D facial dataset constructed from high resolution scans [14] represented in the 2D UV domain. Our model is trained from scratch, without the need for a pre-trained foundation model like Stable Diffusion. To condition our model we train one set of cross-attention layers for each type of conditioning input, following IP-Adapter [13]. First, the diffusion model learns to inject FLAME parameters via the original UNet cross-attention layers. We then freeze the diffusion model while training additional sets of cross-attention layers (e.g. one for artist sketches, 2D landmarks, portrait photos, etc). Our FLAME-conditioned model allows us to re-interpret FLAME-parameterized faces in a generative sense - providing a space of high resolution stochastic variations on top of the traditional low resolution FLAME model.
@@
Our method allows for fast and user-friendly creation of 3D digital character faces with expressions, generated with a consistent mesh topology, and controlled by the input mode preferred by the user, all within a single model. We demonstrate several applications of our method including sketch-based 3D face modeling, geometry from 2D facial landmarks, Canny edges, or portrait photos, text-to-3D facial geometry, and finally, extending the FLAME model space by allowing stochastic diffusion sampling conditioned on the same semantic FLAME parameters (Fig. 1). In summary, we make the following contributions:
1.
We present a new method for 3D face geometry generation from 6 different types of conditionings (prompts) within a single model.
2.
We propose a comprehensive solution for training such a method from scratch, with 3D geometry data augmentations and by representing 3D geometry as position maps to better fit existing diffusion pipelines.
3.
We show that our method supports face generation with expressions, sketch-based editing for 3D face design, stochastic variations of details conditioned on low resolution FLAME faces, generalization to in-the-wild data and dynamic face generation from videos.
2 Related Work
@@
2.1 3D Face Geometry Generation
@@
Recent work uses diffusion models to control the generation of novel 3D face geometry. ShapeFusion [15] generates face geometry by running the diffusion process directly on the mesh input vertices. It allows unconditional and conditional face geometry generation and supports various editing operations on a given mesh, based on selected vertices (anchor points). However, it does not support conditioning signals beyond vertices. 4DFM [16] trains an unconditional diffusion model on a set of sparse 3D landmarks for facial expression generation. It can generate dynamic facial expression sequences based on 3D landmarks by retargeting the landmarks to a mesh after the diffusion process. While they support conditioning with different signals such as expression labels and text, they achieve control via classifier-guidance [17] which requires training additional classifiers on noisy data.
@@
Other methods focus on 3D face or head avatars, which can generate 3D geometry and texture. Rodin [18] can generate triplane-based head geometry with text or image conditioning, but the resulting geometry is extracted with Marching Cubes [19] and thus not in the same topology across generations. HeadArtist [6], HeadSculpt [7] and HumanNorm [20] use Deep Marching Tetrahedra [21], text-prompts and a Score Distillation Loss (SDS) [22] to generate high-quality human heads. However, the topology of the extracted geometry differs across samples and generating a single geometry sample takes almost one hour on a single 3090 GPU. DreamFace [4], FaceG2E [5] and Bergman et al. [23] propose 3D Morphable Model (3DMM)-based [24] pipelines that generate topology-consistent 3D face geometry and textures from text. The geometry is created by optimizing 3DMM parameters using a SDS loss. During the optimization, the SDS loss uses the feedback from a pre-trained text-to-image latent diffusion model to update the 3DMM parameters given a geometry render. DreamFace [4] and FaceG2E [5] focus on optimizing 3DMM identity parameters and have therefore difficulties in directly generating facial expressions from text prompts. However, DreamFace does support generating faces with expressions from image prompts and FaceG2E results can be imported into the CG pipeline where facial expressions can be added in a separate step after the text-based generation. In general, SDS-based methods rely on the Stable Diffusion (SD) [9] prior which was pre-trained on billions of images [25] to guide their generations. SDS optimization is usually much slower in runtime compared to standard diffusion sampling, as it must backpropagate gradients from a diffusion model to a 3D model via a differentiable renderer for many optimization steps. Describe3D [26] can generate 3D face geometry from text-prompts without diffusion, by mapping CLIP text embeddings to 3DMM shape parameters. However, it does not support different facial expressions.
@@
In our work, we generate controllable 3D face geometry in a single common topology from several different conditioning modes. Our method natively supports facial expressions and does not rely on SDS optimization, classifier-guidance or the SD prior.
2.2 Multimodal Conditional Image Generation
@@
Image generation with diffusion models can be controlled with conditioning modes (prompt types) that are different from text such as sketches [27], Canny edges [28], RGB images [13], expression parameters [29] or face shape [30, 31]. To control pre-trained diffusion models with new modes, ControlNet [11] introduces a trainable copy of the diffusion model’s UNet encoding blocks, which take the new conditioning as input. The output of the copied model is added to the skip-connections of the frozen pre-trained diffusion model. A separate trainable copy of the UNet encoding blocks (361M parameters) is created per conditioning mode. T2I-Adapter [12] aligns the internal knowledge of a pre-trained diffusion model with new control modes by proposing a small adapter network that achieves control similar to ControlNet, while requiring less parameters (77M). IP-Adapter [13] injects each conditioning via separate cross-attention layers [32] while requiring even less parameters (22M). It introduces new cross-attention layers whose outputs are added to those of the original UNet. Ye et al. [13] show that the diffusion model follows the added conditioning signal closely, when it comes through the newly trained cross-attention layers. In general, adapters can add control with new conditioning modes even long after the training of the underlying base diffusion model is concluded. They can also add new conditioning modes for which only limited paired training data is available, because the underlying diffusion model is frozen, avoiding issues such as catastrophic forgetting [33, 11]. While pre-trained text-to-image diffusion models understand the RGB image domain and can generate images conditioned on a large variety of input modalities, the domain gap to 3D face geometry is large. Therefore, many related works use the rather slow SDS optimization procedure to lift faces into 3D. To allow for faster inference-time sampling, we train our diffusion model from scratch using ground truth 3D face data and geometry data augmentations. We represent 3D face geometry in the 2D UV domain, which enables us to train a 2D diffusion model that can incorporate new conditioning modes using IP-Adapters.
3 Multimodal 3D Face Geometry Generation
@@
3.1 Dataset and Geometry Representation
@@
To represent our face geometries, we generate a novel dataset based on the 3D scan data acquired by Chandran et al. [14], where all faces are stabilized and in topological correspondence. In total, we use 323 identities in our training dataset, where each identity shows 24 different facial expressions (7752 examples). The full dataset contains 341 identities and we randomly choose around 5% (all expressions of 18 identities) as a validation set. We subtract a template face shape from all faces in the dataset and thus represent each individual face as a delta from the template face. The computed delta representation reduces artifacts in the generated 3D face geometry, when compared to the full face representation. We transform each delta face into a vertex delta position map in UV space, which is suitable for being processed by neural networks [34, 35, 36]. This representation records the x, y, z coordinates of the face geometry within a 3-channel image, similar to traditional color texture maps, but instead of an RGB value at each pixel we store the x, y, z delta values. To improve generalization, we augment our existing geometry training data by synthetic identities which we generate via identity interpolation (50k examples) and by mixing face parts of different identities together [37] (150k examples). Adding the augmentations during LDM training improves the generalization to novel identities when conditioning on the FLAME parameters. We kindly refer to our supplementary material for ablation studies on the geometry representation and the use of geometry data augmentations. Combining original and augmented data leads to a total training dataset size of 207752 examples. Thus our training dataset is slightly larger in size compared to related 2D image diffusion models that specialize on human faces (e.g. 30K images for the CelebA-HQ [38] face LDM [9], or the 70K images for the Diffusion Autoencoder in Preechakul et al. [39]), but still much smaller than the datasets required to train general foundation diffusion models that can represent various objects beyond faces. According to Kadkhodaie et al. [40] diffusion models trained on around 100K samples provide evidence of strong generalization in the face domain. We use the parameter space of a common 3D morphable face model (FLAME [3]) as a base conditioning because we can fit FLAME to the scan data and to the augmented data and thus generate a large dataset of paired geometry-FLAME parameter data. Additionally, we create paired training data for several conditioning modes that only have limited paired data available. For example, portrait photos are only available for the scans from the dataset of Chandran et al. [14], but not for the augmentation data. However, it is possible to inject new modalities with limited paired data by training new cross attention layers while keeping the LDM frozen as shown by Ye et al. [13] (and described in Section 3.4).
3.2 Variational Autoencoder
@@
To reduce the computational requirements for the diffusion model, we downsample our UV position map data by a factor of four into a latent space using a variational autoencoder (VAE) [41] consisting of an encoder and a decoder . We train the VAE from scratch following the autoencoder loss function and architecture as it is presented in related work [9, 42]. Specifically, we use the VQ-GAN [42] autoencoder loss:
@@
consists of a pixel-wise L1 loss and a LPIPS [43] perceptual loss. It compares the input UV position maps to the reconstructions through the VAE. evaluates inputs and reconstructions with a patch-based discriminator [44] and employs a codebook loss which serves as a latent space regularizer. Please refer to Esser et al. [42] for more details.
3.3 Latent Diffusion Model
@@
Next, we train a latent diffusion model (LDM) [9], that learns to generate latent UV position maps . To train the LDM, a forward diffusion process is defined as a Markov chain, which noises the latents following a fixed noise schedule of uniformly sampled timesteps. At the last time step , the distribution is Gaussian. We can directly sample at an arbitrary timestep by:
@@
We learn to predict the noise that was added to a noisy latent image following Ho et al. [45]:
where is the timestep, is a FLAME parameter conditioning and is the UNet [46] neural network with parameters . The FLAME conditioning is obtained by fitting FLAME to the face geometry encoded in and mapping it through a MLP .
3.4 Multimodal Conditional Generation
To control the generations with additional conditioning modes (beyond the FLAME parameters ), we train different sets of cross-attention layers, following IP-Adapter [13]. The LDM itself is kept frozen. In this way, we can integrate novel conditioning modes post-LDM training even with limited paired mode-geometry data. We train one set for each of the following conditioning modes: sketches , Canny edges , 2D landmarks , portrait photos and text . The output of the new cross-attention layers is added to the outputs of the existing LDM cross-attention layers, thereby injecting the new conditioning signal into the generation process:
@@ 3.4 Multimodal Conditional Generation
Prior to passing each of the above-mentioned conditioning modes to the cross-attention layers, we pass each of them through CLIP [48] and extract a 768 dimensional global CLIP feature vector, which serves as our conditioning representation. Following IP-Adapter, we train a small projection network consisting of one linear layer and layer normalization, designed to project the CLIP feature vector into several extra context tokens before injecting it into the cross-attention layers. We use 16 tokens for each of our conditionings to allow for meaningful attention computation.
@@
At inference time, we can control the 3D face geometry generation with any of the modes using the respective set of cross-attention layers and classifier-free guidance [49]. The strength of the conditioning signal can be increased by increasing the hyperparameter :
@@
To condition only on a newly added mode or to generate geometry unconditionally, the FLAME conditioning is set to its null embedding. Additionally, for unconditional generation, the new cross-attention in Eq. 4, which feeds to the diffusion model, is simply not added to the original attention. Our full method pipeline is visualized in Fig. 2. For more implementation details, please refer to our supplementary material.
4 Results